OpenStudy (anonymous):

simplify please help me to simplify this rational equation x^2-2^2/x+2

OpenStudy (anonymous):

I suggest factorising the numerator (tip: they're both squares)

OpenStudy (anonymous):

(x^2-2^2)/(x+2)=[(x-2)(x+2)]/(x+2)=(x-2)